Status Report

Hubble Science Operations Deferred While Engineers Examine New Issues

By SpaceRef Editor
October 17, 2008
Filed under , ,
Hubble Science Operations Deferred While Engineers Examine New Issues

Activation of Hubble Space Telescope science instruments and resumption of science observations has been suspended following two anomalies seen in systems onboard the telescope yesterday.

The first event occurred at approximately 1:40 PM on Thursday when the Advanced Camera for Surveys (ACS) suspended operations due to indication that one power source in the Solar Blind Camera’s low voltage power supply failed to turn on. This was the first turn on of the Solar Blind Camera since its last use prior to the safing of HST’s NASA Standard Spacecraft Computer, version 1 (NSSC-1) on 9/27. (The NSSC-1 is Hubble’s science computer.) The Solar Blind Channel is the only part of ACS in use since a power failure in January 2007 halted most ACS science observations.

An initial status briefing to discuss the ACS issue began at 5 p.m. on Thursday. During the briefing, at 5:14 PM, a still-undiagnosed problem affecting the Science Instrument Command and Data Handling (SI C&DH) system caused the NSSC-1 to stop issuing its keep-alive signal. The Hubble spacecraft computer (HST486) detected the cessation of the keep-alive signal as a failure of its Processor Interface Table (PIT) toggle test. The absence of the keep-alive signal for twenty seconds told the HST486 to issue commands to safe HST’s science computer and science instruments.

The first step taken in investigating the NSSC-1 safing was to collect a memory dump. This dump was success. The data shows that yesterday’s anomaly is likely different from the one that occurred on Sept. 27.

Initial indications point to a potential problem in the B-side of the Computer Processing Module (CPM-B). The CPM-B had not been used on orbit prior to Wednesday evening’s activation of Side B of the Science Instrument Control & Data Handling (SIC&DH) unit.

The spacecraft computer is continuing to execute command loads sent to it from the Space Telescope Operations Control Center (STOCC) at GSFC, and all of its subsystems are nominal.

The investigation is continuing. Contingency procedures for a potential switch to a hybrid SI C&DH configuration that would use portions of its Side A and portions of its Side B were tested this morning in HST Program’s Vehicle Electrical System Test (VEST) Facility, a high fidelity mock-up of Hubble that resides in a cleanroom at Goddard.

NASA managers are expected to meet around 1:30 p.m. today to assess the telescope’s current status and discuss future actions.

A telecon for news media representatives is tentatively planned for 3 p.m. today. A media advisory with details was issued by the NASA Headquarters newsroom earlier today.

SpaceRef staff editor.